OpenStudy (anonymous):

Solve the quadratic equation 3x2 + 5x + 2 = 0 by factoring

OpenStudy (anonymous):

3x^2 +2x+3x +2 now take common factors from both

OpenStudy (anonymous):

3x^2+(3+2)x+2=0 3x^2+3x+2x+2=0 3x(x+1)+2(x+1)=0 (3x+2)((x+1)=0 x=-2/3 or x=-1

OpenStudy (anonymous):

dear pura question solve nahi kiya karte @madrockz

OpenStudy (anonymous):

agr maine pura solution de diya toh kya galat hai?jisne ques post kiya hai wo solutiion dekhke steps smjh v skta hai naa

OpenStudy (anonymous):

moderaters id block kar dete h... teen baar mujhe bol chuke h... ;)

OpenStudy (anonymous):

A. the set of negative one and negative two thirds B. the set of one and negative two thirds C. {2, 5} D. {-2, 5}

OpenStudy (anonymous):

????

OpenStudy (anonymous):

A is the ans

OpenStudy (anonymous):

ohh ok thank you :)